A rope tied vertically to the ceiling is pulled to the right by a force perpendicular (90°) of the rope. The rope is tied to a wall with the rope forming a 120° angle counter-clockwise from the vertical rope.


Find the tension of the rope is F = 100 N


1
Expert's answer
2021-04-01T18:34:20-0400

Given,

When the rope is at the right angle, then tension in the rope "(T)=0"

Let the tension in the string becomes T, when the angle becomes "120^\circ"



let the length of the string be l, now applying the conservation of energy

"mg\times l=\frac{1}{2}mv^2 +mgl\cos30"


"2(l-l\cos30 )= v^2"


Hence, net tension in the string


"T = 100 \cos 30 + \frac{mv^2}{l}"


"=50\sqrt(3)+2m(1-\frac{\sqrt{3}}{2})"